OpenStudy (vortish):

the circumference of a circle is 16 prsquard find its radius and area

OpenStudy (anonymous):

what is prsquard?

OpenStudy (vortish):

3.14

OpenStudy (anonymous):

so circumference of a circle is 16pi?

OpenStudy (anonymous):

assuming that...C = 2piR 16pi = 2piR 8 = Radius A= piR^2 you can do the rest

OpenStudy (vortish):

yes

OpenStudy (anonymous):

did you get this?

OpenStudy (vortish):

no I M stuck where sheng left off \

OpenStudy (vortish):

do you multipy 8 by 3.14 then square

OpenStudy (anonymous):

But you did follow how he got the radius? Circumference = 16 * pi (given in the problem) = 2 pi * radius So 16 * pi = 2 * pi * radius Next, divide both sides by 2*pi: (16 * pi) / (2 * pi) = (2 * pi * radius) / (2 * pi) which simplifies to 8 = radius Area is equal to pi * radius squared, or pi * radius * radius So that would be pi * 8 * 8, or 64 * pi If you have been told to estimate pi as 3.14, that's fine... then you multiply 64 by 3.14 to find the area.

OpenStudy (anonymous):

Shorter answer to your last question, no, you first square the radius, THEN multiply by 3.14

OpenStudy (anonymous):

It's like:\[\pi \times (radius ^{2})\]

OpenStudy (vortish):

ok so if i take 8 times 8 times 3.14 I get 64 times 3.14 which comes out to 200.96

OpenStudy (anonymous):

ok, sounds good to me :) Remember, in case it matters, that the units for area are going to be "square" something... square centimeters or inches or miles or whatever. Length is just like meters or inches, but area is square units.

OpenStudy (vortish):

i typed that in and it was wrong for the area its inches squared

OpenStudy (anonymous):

are you sure you are supposed to use 3.14 for pi?

OpenStudy (anonymous):

Also, your initial post of the question had typos... might be good to double check the question you're seeing online and paste it again or take a screenshot or something... But if you're sure it says "circumference = 16 pi", I will believe it :)

OpenStudy (vortish):

its has that  in it

OpenStudy (vortish):

is in the problem

OpenStudy (anonymous):

it looks like you typed in "64" in the blank for area. The page's instructions say to type in an exact answer in terms of "pi". So don't do that last step where you made "pi" equal to 3.14... just leave it as "pi" So your area would be "pi" x "8 squared", or 64 pi. Type 64 in that blank, then use the pi symbol button over on the left to add the pi symbol in after the 64 in the blank.

OpenStudy (vortish):

k thanks